If A and B be events in a sample space such that P(A)=0.3 , P(B)=0.6 and P(AUB)=0.8 , find P(A' and B') .           (A' means not A)

Grade:

1 Answers

Samyak Jain
5 years ago
P(A) = 0.3, P(B) = 0.6, P(A \small \bigcup B) = 0.8
Suppose S denotes the sample space. Probability of sample space is always 1
i.e. P(S) = 1
P(A’ and B’) = P(A’ \small \bigcap B’) = P(S) – P(A \small \bigcup B) = 1 – 0.8
P(A’ and B’) = 0.2